OpenStudy (bananas1234):

Solve for c. A = 1/4(b + c)

OpenStudy (hockeychick23):

I'm assuming you just need to simplify so you start by taking your equation: A = 1/4(b + c) and multiplying it by 4 (A= 1/4b+1/4c)4 This leaves you with: 4a= b+c Next you subtract b from both sides 4a-b=b-b+c This leaves you with 4a-b= c

OpenStudy (bananas1234):

Thank you!

OpenStudy (hockeychick23):

anytime :)